Monday, Mar. 07: SWING Shift
From Hall A Wiki
Go back to Daily Run Plan Daily Run Plans
- Production data (Kin48_3)
- Continue production mode ***
last update Mar 06
- Polarized beam at 15 uA on LH2 15 cm through Compton chicane !
- DVCS is running kin_48_3 [here], as off 03-06 5 pm we have accumulate 2.14 C out of our goal of 4.7 C.
- LHRS is partially detuned: [here]
- LHRS=2.92GeV, Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.
- beam locks: no position lock, energy lock in Arc 2: [here]
- the straw chambers in the LHRS are ON !
- DVCS daq configuration is [this one], we are measuring about 2% DT at 15uA.
- Take Compton run every 2 h
- After beam lost for considerable time (like beam study, pass change,...)
- Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++
Please check more info on the white board in the Counting House ~!!